Circular No. 8198

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003hw Further to IAUC 8197, J. Graham and W. Li report the LOTOSS

discovery of an apparent supernova (mag ∼ 18.3) on unfiltered KAIT images taken on Sept. 8.5 and 10.5 UT. SN 2003hw is located at α = 3^{h}01^{m}50^{s}.02, δ = +35^{o}44′36′′.4 (equinox 2000.0), which is 2′′.9 east and 2′′.1 north of the nucleus of a galaxy that itself is 49′′ west and 35′′ north of the nucleus of UGC 2491. A KAIT image taken on Feb. 19.2 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2003hv IN NGC 1201 A. Dressler, M. Phillips, N. Morrell, and M. Hamuy, Carnegie Observa-

tories, report that a spectrum (range 390–1007 nm) of SN 2003hv (cf. IAUC 8197), obtained at Las Campanas Observatory on Sept. 10.36 UT with the Baade 6.5-m telescope (+ IMACS), shows it to be a type-Ia supernova. The spectrum bears strong resemblance to that of SN 1994D taken two days after maximum. The minimum of the Si II 635.5-nm line yields an expansion velocity of −10800 km/s (adopting the NED host-galaxy recession velocity of 1671 km/s).

SUPERNOVA 2003ho IN ESO 235-G58 M. Hamuy and M. Roth, Carnegie Observatories, report that a spec-

trum (range 360–900 nm) of SN 2003ho (cf. IAUC 8186), obtained on Sept. 7.04 UT with the Clay 6.5-m telescope (+ LDSS-2) at Las Campanas, shows that this is a normal type-II supernova with a red continuum and broad Balmer P-Cyg lines. The minimum of the Hβ absorption yields an expansion velocity of −5420 km/s (assuming the NED recession velocity of 4265 km/s for the host galaxy), which suggests that the supernova is probably one month old.

COMET C/2002 T7 (LINEAR) Total visual magnitude estimates: Aug. 2.03 UT, 23.09, 13.0 (K. Hor-

noch, Lelekovice, Czech Republic, 0.35-m reflector); 26.78, 12.2 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); 27.14, 12.9 (J. J. Gonzalez, Leon, Spain, 0.20-m reflector); 28.98, 12.2 (V. S. Nevski, Vitebsk, Belarus, 0.30-m reflector); Sept. 4.77, 12.7 (S. Yoshida, Gunma, Japan, 0.25-m reflector); 6.96, 12.3 (S. E. Shurpakov, Baran, Belarus, 0.41-m reflector).

2003 September 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green